For the given sequence 1/n, 1/2n, 1/4n,…  Find the 16th term. Find sum of first sixteen terms where n is 156 


1
Expert's answer
2021-07-16T11:17:47-0400

"{a_1} = \frac{1}{n} = \frac{1}{{{2^0} \cdot n}}"

"{a_2} = \frac{1}{{2n}} = \frac{1}{{{2^1} \cdot n}}"

"{a_3} = \frac{1}{{4n}} = \frac{1}{{{2^2} \cdot n}}"

Then

"{a_k} = \frac{1}{{{2^{k - 1}} \cdot n}}"

So

"{a_{16}} = \frac{1}{{{2^{15}} \cdot n}} = \frac{1}{{{\rm{32768}}n}}"

We find the required sum as the sum of the first 16 terms of the geometric progression with the first term "{a_1} = \frac{1}{n}" and the denominator "q = \frac{1}{2}" :

"{S_{16}} = \frac{{{a_1}\left( {1 - {q^{16}}} \right)}}{{1 - q}} = \frac{{\frac{1}{n}\left( {1 - {{\left( {\frac{1}{2}} \right)}^{16}}} \right)}}{{1 - \frac{1}{2}}} = \frac{{65535}}{{32768n}}"

If "n=156" then

"{S_{16}} = \frac{{65535}}{{32768 \cdot 156}} = \frac{{21845}}{{1703936}}"

Answer: "{a_{16}} = \frac{1}{{{\rm{32768}}n}}" , "{S_{16}} = \frac{{21845}}{{1703936}}"